OpenStudy (anonymous):

lim x->infinity [ (a^x)/{(a^x)+1} ]

OpenStudy (anonymous):

calculus 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

it is also given a>0.

OpenStudy (anonymous):

Multiply top and bottom by \[\frac{1}{a^x}\]

OpenStudy (anonymous):

This is a rational functions problem you might have seen in precalc. Since the degree of the numerator will always be the same as the degree of the denominator as x --> infinity, the function goes to the ratio of the coefficients of the highest degree term of the top and bottom of the fraction.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

which is 1

OpenStudy (anonymous):

\[\frac{1}{1+ \frac{1}{a^x}}\]

OpenStudy (anonymous):

the second term in bottom goes to zero

OpenStudy (anonymous):

we can't say anything about a^x ...because we don't know about "a".(other than only a>0)

OpenStudy (anonymous):

if suppose a=1 then what will happen......i think we will get 1/2